# Feedback-driven recurrent quantum neural network universality

### v1

@@ 2.2 RQNN approximation of state-space maps and their derivatives
2.3 Recurrent QNN approximation bounds for state-space filters
The results in the previous section show that the family of RQNNs that were introduced in (3) is capable of approximating arbitrarily well the very general class of continuously differentiable state-space maps with bounded Fourier transform, together with their derivatives. These approximations hold with respect to both the $L^{2}$ norm (Proposition 2.2 and Corollary 2.3) and the $L^{\infty}$ norm on compacta (Proposition 2.5 and Corollary 2.6). As in the internal approximation approach introduced in [GO18a, Theorem 3.1 (iii)], we will use the uniform RQNN approximation results for the state maps to conclude similar uniform approximation results for the corresponding filters under additional hypotheses that guarantee that those exist.
Consider a state-space system
$\bm{x}_{t}=F(\bm{x}_{t-1},\bm{z}_{t}),\quad t\in\mathbb{Z}_{-},$ (23)
with state process $(\bm{x}_{t})_{t\in\mathbb{Z}_{-}}$ valued in $\mathbb{R}^{N}$, input process $(\bm{z}_{t})_{t\in\mathbb{Z}_{-}}$ valued in $\mathbb{R}^{d}$ and $F\colon\mathbb{R}^{N}\times\mathbb{R}^{d}\to\mathbb{R}^{N}$. We work under the assumption that $F$ satisfies Barron-type integrability conditions [Bar92], [Bar93], [BK18] and $F$ is contractive. Then, see, e.g., Proposition 1 and Remark 2 in [GGO20], it follows that, for any compact subset $D_{d}\subset\mathbb{R}^{d}$, the associated filter $U^{F}\colon(D_{d})^{\mathbb{Z}_{-}}\to(B_{N})^{\mathbb{Z}_{-}}$ induced by the restriction of $F$ to $B_{N}\times D_{d}$ is well-defined and continuous.
Our next result shows that among the RQNNs that we discussed in Proposition 2.5 there exist systems that have the echo state property and hence have a filter associated. More importantly, those filters can be used to uniformly approximate any of the filters corresponding to the general systems introduced above in (23) as long as they satisfy a Barron-type integrability condition and are sufficiently contractive.
Proposition 2.7.
Suppose $F$ in (23) is continuously differentiable with $\|\nabla_{\bm{x}}F_{j}(\bm{x},\bm{z})\|\leq\frac{\lambda}{\sqrt{N}}$ for all $\bm{x}\in\mathbb{R}^{N},\bm{z}\in D_{d}$, $j\in\{1,\ldots,N\}$ for some $\lambda\in(0,1)$ and, moreover, $F$ satisfies $F_{j}\in\mathcal{F}_{R}$, $\partial_{i}F_{j}\in\mathcal{F}$ and $\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\|\bm{\xi}\|^{4}|\widehat{F_{j}}(\bm{\xi})|\mathrm{d}\bm{\xi}<\infty$ for $j=1,\ldots,N$. Denote by $U^{F}\colon(D_{d})^{\mathbb{Z}_{-}}\to(B_{N})^{\mathbb{Z}_{-}}$ the filter associated to (23). Then for any $n\in\mathbb{N}$ with $n>n_{0}$ there exists $\bm{\theta}\in\bm{\Theta}$ such that the system (4) has the echo state property and the associated filter $\bar{U}\colon(D_{d})^{\mathbb{Z}_{-}}\to(\mathbb{R}^{N})^{\mathbb{Z}_{-}}$ satisfies
@@ 2.3 Recurrent QNN approximation bounds for state-space filters
Choose $M$ such that $B_{N}\times D_{d}\subset[-M,M]^{N+d}$ and $[-R,R]^{N}\times D_{d}\subset[-M,M]^{N+d}$. Firstly, our hypotheses on $F$ guarantee that $F$ satisfies the hypotheses of Proposition 2.5. Hence, there exists $\bm{\theta}\in\bm{\Theta}$ such that for any $j\in\{1,\ldots,N\}$,
@@
$\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})\|$ $\leq\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})-\nabla_{\bm{x}}F_{j}(\bm{x},\bm{z})\|+\|\nabla_{\bm{x}}F_{j}(\bm{x},\bm{z})\|$ (26)
$\leq\left(\sum_{i=1}^{N}|\partial_{i}\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})-\partial_{i}F_{j}(\bm{x},\bm{z})|^{2}\right)^{1/2}+\frac{\lambda}{\sqrt{N}}$
$\leq\sqrt{N}\frac{\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\frac{\lambda}{\sqrt{N}}.$
Therefore, using the mean-value theorem, we obtain for all $\bm{x}\in[-M,M]^{N},\bm{z}\in D_{d}$ that
$\|\bar{F}^{n,\bm{\theta}}_{R}(\bm{x}^{1},\bm{z})-\bar{F}^{n,\bm{\theta}}_{R}(\bm{x}^{2},\bm{z})\|^{2}$ $=\sum_{j=1}^{N}(\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x}^{1},\bm{z})-\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x}^{2},\bm{z}))^{2}$
$\leq\sum_{j=1}^{N}\|\bm{x}^{1}-\bm{x}^{2}\|^{2}\max_{\bm{x}\in[-M,M]^{N}}\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})\|^{2}$
$\leq N\left(\sqrt{N}\frac{\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\frac{\lambda}{\sqrt{N}}\right)^{2}\|\bm{x}^{1}-\bm{x}^{2}\|^{2}.$
In particular, for $n$ satisfying $N^{2}\frac{(\max_{j=1,\ldots,N}C_{j}^{\infty})^{2}}{(1-\lambda)^{2}}<n$ we obtain that $\bar{F}^{n,\bm{\theta}}_{R}\colon B_{R}\times D_{d}\to B_{R}$, with $B_{R}=\{\bm{x}\in\mathbb{R}^{N}\colon\|\bm{x}\|\leq R\sqrt{N}\}$, is contractive in the first argument, hence the system (4) has the echo state property by [GGO20, Proposition 1].
By the mean value theorem, the assumption $\|\nabla_{\bm{x}}F_{j}(\bm{x},\bm{z})\|\leq\frac{\lambda}{\sqrt{N}}$ guarantees that $F(\cdot,\bm{z})$ is $\lambda$-contractive for any $\bm{z}\in D_{d}$. Hence, we may estimate
@@
2.4 Universality
with $\bm{\theta}=(\bm{\theta}^{1},\ldots,\bm{\theta}^{N})\in\bm{\Theta}^{N}$ and $P_{1},\ldots,P_{N}\in\mathbb{R}^{N\times N}$ linear preprocessing maps. Our modified recurrent quantum neural network is then defined by the state-space system associated to the state map $\tilde{F}^{n,\bm{\theta}}_{R}$
The next Lemma shows that adding linear preprocessing maps to reservoir equations can lead to the echo state property without contraction assumptions. This approach was first introduced in [GO20] and subsequently used, e.g., in [GGO23, GO21].
Lemma 2.8.
$\tilde{F}_{j}(\bm{x},\bm{z})=g_{j}(P_{j}\bm{x},\bm{z})$ (31)
@@ 2.4 Universality
Notice that Lemma 2.8 provides the echo state property by imposing a finite memory of $K-1$ time steps on the reservoir. Let $D_{d}\subset\mathbb{R}^{d}$, $B_{m}\subset\mathbb{R}^{m}$ be compact. For a readout $W\in\mathbb{R}^{m\times N}$, denote
$\mathbf{y}_{t}=W\bm{x}_{t}$ (33)
the output process associated to the recurrent QNNs (4) and (30).
Theorem 2.9.
@@
Let $H_{U}\colon(D_{d})^{\mathbb{Z}_{-}}\to(B_{m})^{\mathbb{Z}_{-}}$ be the functional associated to the filter $U$. Then, as in the proof of [GO21, Theorem 2.1], there exists $K\in\mathbb{N}$ and a continuous function $\bar{G}\colon(D_{d})^{dK}\to B_{m}$ such that
@@
Moreover, e.g., by the argument in [GJ25, Corollary 4], there exists a function $G\in C^{\infty}_{c}((\mathbb{R}^{d})^{K},B_{m})$ which satisfies
@@
Next, choose $N=(K-1)d+m$ and consider the recurrent QNN introduced in (4). Denote
@@
Without loss of generality we may choose $R=R_{1}=\ldots=R_{K}$, since we can always replace $R_{k}$ by $\max(R_{k},R_{k+1})$ (and hence ultimately replace $R_{1},\ldots,R_{K}$ by $R$) and absorb the change in an adjusted choice of parameters $\gamma^{i,j}$ (see representation (7)). Moreover, by a similar reasoning we may assume without loss of generality that $n=n_{1}=\ldots=n_{K}$. Indeed, otherwise we may again choose $n$ to be the maximum of $n_{1},\ldots,n_{K}$, replace $n_{1},\ldots,n_{K}$ by $n$ and recover the same functions (7) by setting surplus terms $i>n_{k}$ to 00 by appropriate choice of $\gamma^{i,j}$. The extra factor $\frac{n}{n_{k}}$, in turn, can be absorbed by modifying the choice of $R$.
@@
Assume that the bound holds for a fixed $k\in\{K,\ldots,3\}$, then for $k-1$ we estimate (with the notation $f_{k-1}=\bar{F}^{n,\bm{\theta}}_{R,l_{k-2}+1:l_{k-1}}$)
@@
3 Conclusions
Approximation bounds and universality properties are part of the theoretical cornerstone of machine learning models. While some studies have addressed the question of universality for QRC models, the combination of the two had not previously been explored in the context of recurrent QNNs. In this paper, we derived approximation bounds and universality statements for recurrent QNNs based on the circuit implementation presented in [GJ25], which is compatible with hardware deployment and whose implementation with Rydberg atoms has been already discussed in [APB+24]. This circuit uses a uniformly controlled quantum gate to apply multi-controlled rotations to a set of control and target qubits, and it has been recently shown that it can be efficiently implemented [ZB24, SAAdS24, ZB25].
To extend the results of [GJ25] to the recurrent case, we first derived approximation bounds for the static version of the QNN and its derivatives. This yields uniform approximation bounds of state-maps that satisfy a Barron-type condition (Proposition 2.5) and a universality statement (Corollary 2.6) for continuously differentiable state-equations. These results are used in Proposition 2.7 to provide filter approximation bounds that show that RQNNs are able to uniformly approximate the filters induced by any contracting Barron-type state-space system. Finally, Theorem 2.9 extends this universality property to the much larger category of arbitrary fading memory, causal, and time-invariant filters. In this last result, neither Barron-type integrability nor contractivity conditions are needed for the target filter. Instead, we modify the approximating recurrent QNN with preprocessing matrices to ensure the echo state property. Note that all these results apply, strictly speaking, exclusively to variational systems for which all parameters are trained and are only tentative for quantum reservoir systems in which some parameters in the recurrent layer are randomly generated. While the variational approach seeks the optimal parameters for the quantum circuit and output layer, the reservoir approach is simpler, usually involving random sampling of the circuit parameters and tuning of only the output layer. Which strategy is best in terms of speed and accuracy will depend on the number of blocks $n$ of the circuit, the intrinsic noise of the hardware, and the target task. Future research will focus on implementing and comparing the variational and reservoir approaches.
This work paves the way for extending the theoretical analysis of QRC models beyond the SAS paradigm [MPO23]. It is important to understand in which situations the feedback approach is preferable to other protocols. Questions such as the exponential concentration of observables [SGZ25, XHA+25] and the suitability of QRC models for learning quantum temporal tasks [TN21, Nok23] are fundamental to discerning the conditions that render QRC models more useful than classical machine learning approaches.
Acknowledgments. The authors acknowledge partial financial support from the School of Physical and Mathematical Sciences of the Nanyang Technological University through the SPMS Collaborative Research Award 2023 entitled “Quantum Reservoir Systems for Machine Learning”. RMP acknowledges the QCDI project funded by the Spanish Government. JPO wishes to thank the hospitality of the Donostia International Physics Center and LG and RMP that of the Division of Mathematical Sciences of the Nanyang Technological University, during the academic visits in which some of this work was developed.

### v2

@@ Appendix A Quantum reservoir computing protocols
Appendix C Proofs for Section 4.2
C.1 Proof of Theorem 4.6
@@
Choose $M$ such that $B_{N}\times D_{d}\subset[-M,M]^{N+d}$ and $R,R]^{N}\times D_{d}\subset[-M,M]^{N+d}$. Firstly, our hypotheses on $F$ guarantee that $F$ satisfies the hypotheses of Proposition 4.4. Hence, there exists $\bm{\theta}\in\bm{\Theta}$ such that for any $j\in\{1,\ldots,N\}$,
@@
$\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R}(\bm{x},\bm{z})\|_{2}$ $\leq\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R}(\bm{x},\bm{z})-\nabla_{\bm{x}}F(\bm{x},\bm{z})\|_{2}+\|\nabla_{\bm{x}}F(\bm{x},\bm{z})\|_{2}$ (32)
$\leq\left(\sum_{i,j=1}^{N}|\partial_{i}\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})-\partial_{i}F_{j}(\bm{x},\bm{z})|^{2}\right)^{1/2}+\lambda$
$\leq N\frac{\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\lambda.$
Therefore, using that $\max_{\bm{x}\in[-M,M]^{N}}\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R}(\bm{x},\bm{z})\|_{2}$ is the best Lipschitz-constant for $\bar{F}^{n,\bm{\theta}}_{R}$ on $M,M]^{N}$ for any given $\bm{z}\in D_{d}$, we obtain for all $\bm{x}\in[-M,M]^{N},\bm{z}\in D_{d}$ that
$\|\bar{F}^{n,\bm{\theta}}_{R}(\bm{x}^{1},\bm{z})-\bar{F}^{n,\bm{\theta}}_{R}(\bm{x}^{2},\bm{z})\|^{2}$ $\leq\|\bm{x}^{1}-\bm{x}^{2}\|^{2}\max_{\bm{x}\in[-M,M]^{N}}\|\nabla_{\bm{x}}\bar{F}^{n,\bm{\theta}}_{R}(\bm{x},\bm{z})\|_{2}^{2}$
$\leq\left(N\frac{\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\lambda\right)^{2}\|\bm{x}^{1}-\bm{x}^{2}\|^{2}.$
In particular, for $n$ satisfying $N^{2}\frac{(\max_{j=1,\ldots,N}C_{j}^{\infty})^{2}}{(1-\lambda)^{2}}<n$ we obtain that $\bar{F}^{n,\bm{\theta}}_{R}\colon B_{R}\times D_{d}\to B_{R}$, with $B_{R}=\{\bm{x}\in\mathbb{R}^{N}\colon\|\bm{x}\|\leq R\sqrt{N}\}$, is contractive in the first argument, hence the system (4) has the echo state property by Gonon et al. (2020, Proposition 1).
By the relation between the Lipschitz-constant and the maximal norm of the Jacobian, the assumption $\|\nabla_{\bm{x}}F(\bm{x},\bm{z})\|_{2}\leq\lambda$ guarantees that $F(\cdot,\bm{z})$ is $\lambda$-contractive for any $\bm{z}\in D_{d}$. Hence, we may estimate
@@
C.2 Proof of Lemma 4.7
The proof of Lemma 4.7 is related to the approach introduced in Gonon & Ortega (2020) and subsequently used, e.g., in Gonon et al. (2023); Gonon & Ortega (2021).
@@
C.3 Proof of Theorem 4.8
@@
Let $H_{U}\colon(D_{d})^{\mathbb{Z}_{-}}\to B_{m}$ be the functional associated to the filter $U$. Then, as in the proof of Gonon & Ortega (2021, Theorem 2.1), there exists $K\in\mathbb{N}$ and a continuous function $\bar{G}\colon(D_{d})^{dK}\to B_{m}$ such that
@@
Moreover, e.g., by the argument in Gonon & Jacquier (2025, Corollary 4), there exists a function $G\in C^{\infty}_{c}((\mathbb{R}^{d})^{K},B_{m})$ which satisfies
@@
Next, choose N=(K1)d+mN=(K-1)d+m and consider the recurrent QNN introduced in (4). Denote
@@
Without loss of generality we may choose $R=R_{1}=\ldots=R_{K}$, since we can always replace $R_{k}$ by $\max(R_{k},R_{k+1})$ (and hence ultimately replace $R_{1},\ldots,R_{K}$ by $R$) and absorb the change in an adjusted choice of parameters $\gamma^{i,j}$ (see representation (7)). Moreover, by a similar reasoning we may assume without loss of generality that $n=n_{1}=\ldots=n_{K}$. Indeed, otherwise we may again choose $n$ to be the maximum of $n_{1},\ldots,n_{K}$, replace $n_{1},\ldots,n_{K}$ by $n$ and recover the same functions (7) by setting surplus terms $i>n_{k}$ to 0 by appropriate choice of $\gamma^{i,j}$. The extra factor $\frac{n}{n_{k}}$, in turn, can be absorbed by modifying the choice of $R$.
@@
Assume that the bound holds for a fixed $k\in\{K,\ldots,3\}$, then for k1k-1 we estimate (with the notation $f_{k-1}=\bar{F}^{n,\bm{\theta}}_{R,l_{k-2}+1:l_{k-1}}$)
@@
Appendix D Construction of $\mathtt{V}$
In this appendix we provide further details on the choice of $\mathtt{V}$ appearing in the quantum circuit. Our presentation follows Gonon & Jacquier (2025).
[…]
Example
[…]
In the case $n_{0}=0$, there is an explicit construction of $\psi$ in terms of Hadamard gates acting on the control qubits. Indeed, for $\mathfrak{n}\geq 2$, (Gonon & Jacquier, 2025, Lemma A.2) shows that
[…]
Appendix E Monte Carlo Error
In practice, the empirical sampling error leads to an additional error component of order $1/\sqrt{S}$ for $S$ independent shots, see, e.g., Qi et al. (2023); Liu et al. (2025). Here, we outline how this Monte Carlo error could be taken into account in the present setting.
More specifically, our QNNs in (3) and (4) are defined using probabilities, rather than their Monte Carlo estimates
$\widehat{\mathbb{P}}_{m}^{n,\bm{\theta}}:=\frac{1}{S}\sum_{s=1}^{S}\mathbbm{1}_{\{m,4+m,\ldots,4(n-1)+m\}}(i^{(s)}),$
with $i^{(s)}$ the measured state in the $i$-th shot. To obtain refined bounds incorporating the sampling error, one would proceed as follows. Denote by $\bar{F}^{n,\bm{\theta},S}_{R}$ the RQNN state map with output probabilities estimated by $S$ shots, by $\hat{\bm{x}}^{S}$ the associated state and by $\bar{U}_{S}$ the associated filter.
For the state map itself, the $L^{2}$-error can be directly controlled (as in Gonon & Jacquier (2025)) by
$\mathbb{E}\left[\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\left|\bar{F}^{n,\bm{\theta}}_{R,j}(\bm{x},\bm{z})-\bar{F}^{n,\bm{\theta},S}_{R,j}(\bm{x},\bm{z})\right|^{2}\mu(\mathrm{d}\bm{x},\mathrm{d}\bm{z})\right]^{1/2}$ (46)
$\leq 2R\sum_{i=1}^{2}\left(\int_{\mathbb{R}^{N}\times\mathbb{R}^{d}}\mathbb{E}\left[\left|\mathbb{P}_{i}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})-\widehat{\mathbb{P}}_{i}^{n,\bm{\theta}^{j}}(\bm{x},\bm{z})\right|^{2}\right]\mu(\mathrm{d}\bm{x},\mathrm{d}\bm{z})\right)^{1/2}$
$\leq\frac{4R}{\sqrt{S}},$
using that $\mathbb{E}[|\mathbb{E}[X_{1}]-\frac{1}{S}\sum_{s=1}^{S}X_{s}|^{2}]=\frac{\mathrm{Var}(X_{1})}{S}$ for i.i.d. random variables $X_{1},\ldots,X_{S}$.
For the associated filter, one may proceed as follows. Firstly, (33) in the proof of Theorem 4.6 can be adapted to
$\left\|\bar{U}_{S}(\bm{z})_{t}-U(\bm{z})_{t}\right\|=\left\|\hat{\bm{x}}_{t}^{S}-\bm{x}_{t}\right\|=\left\|\bar{F}^{n,\bm{\theta},S}_{R}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})-F(\bm{x}_{t-1},\bm{z}_{t})\right\|$ (47)
$\leq\left\|F(\bm{x}_{t-1},\bm{z}_{t})-F(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})\right\|+\left\|F(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})-\bar{F}^{n,\bm{\theta},S}_{R}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})\right\|$
$\leq\lambda\left\|\bm{x}_{t-1}-\hat{\bm{x}}_{t-1}^{S}\right\|+\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}-\bar{F}^{n,\bm{\theta}}_{R,j}\right\|_{\infty,M}^{2}\right)^{1/2}+\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta}}_{R,j}-F_{j}\right\|_{\infty,M}^{2}\right)^{1/2}$
$\leq\lambda\left\|\bm{x}_{t-1}-\hat{\bm{x}}_{t-1}^{S}\right\|+\frac{\sqrt{N}\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}-\bar{F}^{n,\bm{\theta}}_{R,j}\right\|_{\infty,M}^{2}\right)^{1/2}.$
The last error term can be bounded as
$\mathbb{E}\left[\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}-\bar{F}^{n,\bm{\theta}}_{R,j}\right\|_{\infty,M}^{2}\right)^{1/2}\right]\leq\left(\sum_{j=1}^{N}\mathbb{E}\left[\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}-\bar{F}^{n,\bm{\theta}}_{R,j}\right\|_{\infty,M}^{2}\right]\right)^{1/2}\leq\frac{C}{\sqrt{S}}$
for a suitable constant $C$ using techniques from statistical learning theory, provided that $\widehat{\mathbb{P}}_{m}^{n,\bm{\theta}}$ is Lipschitz continuous as a function of $(\bm{x},\bm{z})$. Inserting this into (47) and proceeding precisely as in the proof of Theorem 4.6 then yields a bound that incorporates also the sampling error.
Alternatively, as the Lipschitz continuity may be hard to verify, we may obtain an $L^{2}$-bound analogously to Theorem 4.6 as follows. First, using that the shots are independent across evaluations, we may apply (46) to estimate
$\mathbb{E}\left[\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})-\bar{F}^{n,\bm{\theta}}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})\right\|^{2}\right)^{1/2}\right]$
$\quad\leq\left(\sum_{j=1}^{N}\mathbb{E}\left[\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})-\bar{F}^{n,\bm{\theta}}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})\right\|^{2}\right]\right)^{1/2}$
$\quad\leq\sqrt{N}\frac{4R}{\sqrt{S}},$
where the expectations are taken with respect to sampling the probabilities to evaluate $\bar{F}^{n,\bm{\theta},S}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})$.
Next, by proceeding as in (47), we may estimate
$\mathbb{E}[\left\|\bar{U}_{S}(\bm{z})_{t}-U(\bm{z})_{t}\right\|]\leq\lambda\left\|\bm{x}_{t-1}-\hat{\bm{x}}_{t-1}^{S}\right\|+\frac{\sqrt{N}\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}$ (48)
$\qquad+\mathbb{E}\left[\left(\sum_{j=1}^{N}\left\|\bar{F}^{n,\bm{\theta},S}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})-\bar{F}^{n,\bm{\theta}}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})\right\|^{2}\right)^{1/2}\right]$
$\leq\lambda\left\|\bm{x}_{t-1}-\hat{\bm{x}}_{t-1}^{S}\right\|+\frac{\sqrt{N}\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\sqrt{N}\frac{4R}{\sqrt{S}}$
with the expectations again taken with respect to sampling the probabilities to evaluate $\bar{F}^{n,\bm{\theta},S}_{R,j}(\hat{\bm{x}}_{t-1}^{S},\bm{z}_{t})$. In particular, taking expectations also with respect to a random process $\bf Z$} (taking values in $(D_{d})^{\mathbb{Z}_{-}}$) and sampling at each evaluation, the estimate (48) and the same arguments as in the proof of Theorem 4.6 yield the bound
$\sup_{t\in\mathbb{Z}_{-}}\mathbb{E}[\left\|U^{F}({\bf Z})_{t}-\bar{U}_{S}({\bf Z})_{t}\right\|]\leq\frac{1}{1-\lambda}\left(\frac{\sqrt{N}\max_{j=1,\ldots,N}C_{j}^{\infty}}{\sqrt{n}}+\sqrt{N}\frac{4R}{\sqrt{S}}\right).$ (49)